Markson Company had the following results of operations for the past year:

Sales (8,000 units at $20.70) $ 165,600
Variable manufacturing costs $ 88,800
Fixed manufacturing costs 15,700
Variable selling and administrative expenses 14,800
Fixed selling and administrative expenses 20,700 (140,000 )
Operating income $ 25,600


A foreign company whose sales will not affect Markson's market offers to buy 2,000 units at $15.05 per unit. In addition to variable manufacturing costs, selling these units would increase fixed overhead by $1,670 for the purchase of special tools. Markson’s annual productive capacity is 12,000 units. If Markson accepts this additional business, its profits will:

Multiple Choice

  • Decrease by $1,670.

  • Decrease by $5,870.

  • Decrease by $4,900.

  • Increase by $4,200.

  • Increase by $2,530.
Answer #1
Incremental Revenue $          30,100 =2000*15.05
Incremental Variable Manufacturing Costs $        -22,200 =-2000*11.1
Incremental Variable Selling Expenses $           -3,700 =-2000*1.85
Incremental Fixed Costs $           -1,670
Incremental Income $            2,530

Answer is e. Increase by $2,530.